Find the Torque about the rotation produced by constant force of






(6 + 8 − 10 ) 𝑁 on a particle situated at (3 − 4 ) 𝑚.

1
Expert's answer
2021-11-08T17:10:16-0500

Vectors given:


"\vec F=(6\textbf i+8\textbf j-10\textbf k),\\
\vec r=(3\textbf i-4\textbf j+0\textbf k)."

Torque:


"\vec\tau=\vec F\times\vec r,\\
\tau=(8·0-(-10)·(-4),\space(-10)·3-6·0,\space6·(-4)-8·3),\\
\vec\tau=(-32,-30,-48)\text{ N}·\text m."